# Learning with less: label-efficient land cover classification at very high spatial resolution using self-supervised deep learning

### v1

@@ 3 Methodology
Our proposed methodology follows a multi-stage approach to model development that investigated the implications of self-supervised pre-training with the goal of producing an accurate VHSR land cover product over the state of Mississippi using a limited amount of annotated training data (Figure 2). In Section 3.1, we describe a sampling strategy designed to capture the diversity of land cover across Mississippi. Candidate patches were selected and grouped based on their land cover composition, and a subset was sampled for dense visual annotation. A point-based assessment dataset of 25,000 samples was also created for model evaluation and final product accuracy assessment. Section 3.2 describes how BYOL was used to pre-train a ResNet-101 CNN using 377,921 patches of unannotated pre-training data from the area of interest (excluding the already sampled regions). Section 3.3 describes the fine-tuning and evaluation process, where the pre-trained ResNet-101 backbone was integrated into various semantic segmentation architectures and fine-tuned on limited annotated training data using stratified k-fold cross-validation. We evaluated the performance of multiple segmentation architectures (FCN, U-Net, Attention U-Net, DeepLabV3+, UPerNet, PAN) using varying amounts of training data (250, 500, 750 samples) to understand the implications of model architecture and training dataset size on final model performance. Finally, several fine-tuned models were compiled into a model ensemble to produce a land cover map at resolution over the state of Mississippi for the year 2023, with the final data product undergoing a thematic accuracy assessment using the point-based assessment dataset, as described in Section 3.4.
@@ 3.1 Stratified sampling and high-quality land cover training dataset generation
Stratifying the sampling of training data is a common practice in remote sensing to ensure that the training dataset is representative of the land cover distribution of the area of interest, particularly when the area contains rare or underrepresented classes [74, 75, 76]. To ensure representative sampling for k-fold cross-validation, we employed a stratified sampling strategy to select 1,000 candidate training samples for annotation. We overlaid a grid of 256256 pixels (i.e., 256 patches), forming a population of 1,889,606 candidate patches used to support both the unlabeled dataset for pre-training and the labeled dataset for fine-tuning. To curate the fine-tuning dataset, zonal histograms of land cover class distribution were extracted from the 2023 NLCD dataset for each candidate patch to characterize its land cover composition as a type of feature vector. We then applied PCA to reduce the dimensionality of the feature vectors and K-means clustering to form 250 strata based on land cover composition. Finally, we randomly sampled 4 candidate training samples from each stratum, placing each into a separate fold of the k-fold training dataset, resulting in 4 folds totaling 1,000 samples. Figure 3 shows the NLCD land cover across Mississippi, along with the spatial distribution of the selected samples and the class composition within each fold.
@@
A team of trained annotators visually annotated the selected 1,000 samples according to the land cover classification scheme in Table 1. Samples were randomly assigned to each annotator to ensure that each fold contained a mix of contributions from different individuals. After the annotators completed their work, the annotations were reviewed and corrected by another annotator whose task was to ensure that the annotations were consistent and accurate across all samples. Annotators employed the Segment Anything Model 2 (SAM 2) hosted on the Computer Vision Annotation Tool (CVAT) annotation platform to assist with the visual annotation process [77, 78] by proposing initial object masks that annotators could manually refine as needed. Examples of annotated samples are shown in Figure 4.
| Index | Class | |
| 1 | Open Water | |
| 2 | Impervious Structures | |
| 3 | Impervious Surfaces | |
| 4 | Barren Land | |
| 5 | Forest/Woody Vegetation | |
| 6 | Herbaceous/Low Vegetation | |
| 7 | Cultivated Crops | |
| 8 | Unclassified |
@@
3.2 Self-supervised pre-training using BYOL
The pre-training phase of the methodology leverages the large dataset (377,921) of unannotated samples such that a ResNet-101 encoder can be trained to extract relevant features from the imagery that are well-suited for downstream land cover classification tasks. BYOL uses a “self-distillation” approach to learn representations from images using two separate but interdependent networks: an online network containing the image encoder, a projection head, and a prediction head, and a target network that is structurally similar to the online network but lacks the prediction head, initialized using a copy of the online network’s parameters at the start of training [54]. Given an input image, two separate augmented views of the image are created using a set of transformations. As illustrated in Figure 5, each view is then passed through the online and target networks to produce feature vectors. The online network, which processes view , is then trained to predict the feature vector produced by the target network from view . However, the target network is not trained directly; instead, it is updated using an EMA of the online network weights, giving the online network a stable yet constantly updated target to learn from. Ultimately, the image encoder from the online network is retained after training to be transferred to downstream tasks.
@@ 3.2 Self-supervised pre-training using BYOL
We applied BYOL to a ResNet-101 encoder network using the dataset of 377,921 unannotated samples from the state of Mississippi. We tested two configurations of the BYOL pre-training process: one starting with ImageNet pre-trained weights at the beginning of the pre-training process and one starting with randomly initialized weights. The pre-training process was run for 300 epochs, with the learning rate reduced from to 0 over the course of training using a cosine annealing schedule, with a warmup period of 10 epochs to linearly increase the learning rate from 0 to . The AdamW optimizer was used with a batch size of 4096: to achieve this large batch size while working within the memory constraints of a single Nvidia A100 GPU with 80 GB of VRAM, we employed gradient accumulation using a microbatch size of 256. After each epoch, the model was evaluated on the pre-training validation dataset using the BYOL loss function (as during training) to monitor for overfitting.
@@
To evaluate the quality of the learned representations, we used a linear probing approach to determine how well the deep features extracted by the pre-trained networks could be used to classify land cover classes using a simple linear classifier. To accomplish this, the feature maps produced by the ResNet-101 encoder were upsampled to the original input resolution using bilinear interpolation, and a linear classifier was trained on top of these feature maps using the annotated training dataset from Section 3.1. The linear classifier consisted of a single 11 convolutional layer with a softmax activation function to produce class probabilities for each pixel. K-fold cross-validation was employed to minimize the influence of variation in the training dataset while evaluating the performance of this method across different training dataset sizes (); the splitting strategy for each is shown in Figure 6.
@@
A warmup period of 10 epochs was used to linearly increase the learning rate from to at epoch 10, after which the learning rate was decayed by a factor of 0.1 when the validation loss failed to decrease for 10 epochs. After 50 epochs of no improvement in the validation loss, training was halted and the model with the lowest validation loss was selected as the final model for that run. Automatic mixed precision (AMP) training was used to speed up the training stage and reduce memory usage. Due to the limited size of the training datasets, we employed a strong set of image augmentations on-the-fly to ensure that the models did not overfit to the training data, including random horizontal and vertical flips (50% probability each), random 90∘ rotations (0∘, 90∘, 180∘, 270∘; 100%), color jitter (brightness = 0.4, contrast = 0.4, saturation = 0.2, hue = 0.2; 80%), random grayscale conversion (20%), and Gaussian blur (kernel size = 25, ; 100%). Since each cross-validation run used its validation split to tune the optimizer and trigger early stopping, we did not use validation-split metrics for final performance reporting. Rather, the performance of the linear probing models was evaluated using the point-based assessment dataset (25,000 samples), with the results being compared between the two BYOL pre-training configurations (i.e., different weight initialization strategies) as well as a baseline model that used the ResNet-101 encoder with ImageNet pre-trained weights without any BYOL pre-training. Considering the natural imbalance of the land cover classes, we chose the macro F1 score as the primary evaluation metric, as shown in Equation 4, where is the number of classes, is the number of true positives for class , is the number of false positives for class , and is the number of false negatives for class .
@@
Once the ResNet-101 encoders were pre-trained using BYOL, they were used as a backbone for a semantic segmentation architecture that was fine-tuned using the annotated training dataset from Section 3.1. We evaluated 6 different semantic segmentation architectures to determine which framework was best suited for the land cover classification task at hand. We excluded any architectures that rely on transformer-based encoders (e.g., SegFormer [81], SETR [82]) or specialized convolutional backbones (e.g., HRNet [83], BiSeNetV2 [84]) to maintain a controlled comparison between architectures that all use the same ResNet-101 encoder network. Table 2 shows the semantic segmentation architectures that were evaluated in this study, and a brief description of each architecture is provided below.
| Model | Number of Parameters | |
| FCN | 42,528,856 | [85] |
| U-Net | 92,302,024 | [86] |
| Attention U-Net | 83,610,272 | [87] |
| DeepLabV3+ | 58,199,128 | [88] |
| UPerNet | 73,251,400 | [89] |
| PAN | 140,688,970 | [90] |
@@
U-Net is a popular semantic segmentation architecture that was among the early adopters of the encoder-decoder framework for semantic segmentation tasks. Under such a framework, the encoder network extracts hierarchical features from the input image, while the decoder network learns a mapping from these low-dimensional features to an output segmentation map that matches the original input resolution [86]. Notably, U-Net’s decoder network is approximately symmetric to the encoder network, meaning that the feature maps are upsampled in stages that mirror the downsampling stages of the encoder. As such, skip connections can be used to pass feature maps from the encoder to the decoder at corresponding stages: the encoder’s feature maps are concatenated with the decoder’s feature maps at each stage to provide the decoder with relevant, resolution-specific information that otherwise would be lost during the downsampling process. U-Net’s simple yet powerful architecture has made it a popular choice for a wide variety of remote sensing applications [94, 95, 96, 97] Our U-Net implementation differed from the original U-Net architecture in two key ways: first, we opted to use simple bilinear interpolation for upsampling in the decoder rather than transposed convolutions to reduce the presence of checkerboard artifacts in the output segmentation maps (see [98] for more details); second, as we used a ResNet-101 encoder that downsamples the input image by a factor of 4 in the first stage (as opposed to the factor of 2 downsampling in the original U-Net), we modified the decoder to account for this difference by adding an additional upsampling stage to ensure that the output segmentation map matches the input resolution.
@@
DeepLabV3+ is an evolution of the DeepLab family of semantic segmentation architectures that incorporates atrous (dilated) convolutions and spatial pyramid pooling to capture multi-scale context in the feature maps [88]. At the core of DeepLabV3+ is an Atrous Spatial Pyramid Pooling (ASPP) module that aggregates multi-scale context from the final feature maps produced by the encoder. This ASPP module uses parallel atrous convolutions to encode feature maps at multiple scales as opposed to using pooling operations: the use of atrous (i.e., dilated) convolutions allows the network to capture features at multiple scales in a learned manner. A comparatively simple decoder network is utilized that upsamples the output of the ASPP module using bilinear interpolation, concatenates it with low-level feature maps from the encoder, and upsamples the result to produce the final segmentation map. Like U-Net, DeepLabV3+ has seen wide adoption for remote sensing applications [99, 100, 101] DeepLabV3+ was originally designed to use a modified ResNet-101 encoder, and as such minimal effort was necessary to replace the original encoder with our pre-trained ResNet-101 encoder.

### v2

@@ 3 Methodology
Our proposed methodology followed a multi-stage approach to model development that investigated the implications of self-supervised pre-training with the goal of producing an accurate VHSR land cover product over the state of Mississippi using a limited amount of annotated training data (Figure 2). In Section 3.1, we describe a sampling strategy designed to capture the diversity of land cover across Mississippi. Candidate patches were selected and grouped based on their land cover composition, and a subset was sampled for dense visual annotation. A point-based assessment dataset of 25,000 samples was also created for model evaluation and final product accuracy assessment. Section 3.2 describes how BYOL and MoCoV2 were used to pre-train a ResNet-101 CNN using 377,921 patches of unannotated pre-training data from the area of interest (excluding the already sampled regions). Section 3.3 describes the fine-tuning and evaluation process, where the pre-trained ResNet-101 backbone was integrated into various semantic segmentation architectures and fine-tuned on limited annotated training data using stratified k-fold cross-validation. We evaluated the performance of multiple segmentation architectures (FCN, U-Net, Attention U-Net, DeepLabV3+, UPerNet, PAN) using varying amounts of training data (250, 500, 750 samples) to understand the implications of pre-training protocol, model architecture, and training dataset size on final model performance. Several fine-tuned models were compiled into a model ensemble to produce a land cover map at resolution over the state of Mississippi for the year 2023, with the final data product undergoing a thematic accuracy assessment using the point-based assessment dataset, as described in Section 3.4. Finally, in Section 3.5 we describe how we applied our model to 2016 NAIP imagery to create a multi-temporal land cover product and evaluated the generalization capabilities of the final model ensemble through the use of a separate ground truth validation dataset for 2016 using locations sampled in Section 3.1.
@@ 3.1 Stratified sampling and high-quality land cover training dataset generation
Stratifying the sampling of training data is a common practice in remote sensing to ensure that the training dataset is representative of the land cover distribution of the area of interest, particularly when the area contains rare or underrepresented classes [111, 155, 153]. To ensure representative sampling for k-fold cross-validation, we employed a stratified sampling strategy to select 1,000 candidate training samples for annotation. We overlaid a grid of 256256 pixels (i.e., 256 patches), forming a population of 1,889,606 candidate patches used to support both the unlabeled dataset for pre-training and the labeled dataset for fine-tuning. The patch size was chosen following Diakogiannis et al. [27] and Li et al. [68], with 256256 pixels offering sufficient spatial context for supporting high classification accuracy while still remaining computationally efficient [56, 101].
To curate the fine-tuning dataset, zonal histograms of land cover class distribution were extracted from the 2023 NLCD dataset for each candidate patch to characterize its land cover composition as a type of feature vector. We then applied PCA to reduce the dimensionality of the feature vectors and K-means clustering to form 250 strata based on land cover composition. Finally, we randomly sampled 4 candidate training samples from each stratum, placing each into a separate fold of the k-fold training dataset, resulting in 4 folds totaling 1,000 samples. Figure 3 shows the NLCD land cover across Mississippi, along with the spatial distribution of the selected samples and the class composition within each fold. A nearest neighbor analysis revealed a statistically significant spatial clustering effect (), though this clustering effect is mild ().
@@
A team of trained annotators visually annotated the selected 1,000 samples according to the land cover classification scheme in Table 1. Samples were randomly assigned to each annotator to ensure that each fold contained a mix of contributions from different individuals. After the annotators completed their work, all annotations were reviewed and corrected by a single annotator. Only one annotator was used for the quality assurance phase of the annotation process to reduce the effects of inter-annotator disagreement and to enforce a consistent, high quality annotation scheme across the entirety of the ground truth dataset. Annotators employed the Segment Anything Model 2 (SAM 2) hosted on the Computer Vision Annotation Tool (CVAT) annotation platform to assist with the visual annotation process [24, 100] by proposing initial object masks that annotators could manually refine as needed. If SAM 2 was unable to accurately mask an object or feature in the imagery, annotators fell back to manual annotation. Examples of annotated samples are shown in Figure 4.
| Index | Class | Description |
| 1 | Open Water | Rivers, lakes, ponds, and other standing or moving water. Water bodies typically appear dark in CIR imagery due to near-infrared absorption; may appear bluish-green depending on depth and turbidity. |
| 2 | Impervious Structures | Elevated built features (e.g., buildings, sheds). Bright and reflective with sharp edges and shadows. |
| 3 | Impervious Surfaces | Paved areas such as roads and parking lots. Typically bright but may appear dark depending on material. Linear patterns and lack of shadows distinguish from barren land and impervious structures. |
| 4 | Barren Land | Exposed soil, sand, or rock with minimal vegetation. Bright or neutral in CIR imagery, often interleaved with herbaceous areas. |
| 5 | Forest/Woody Vegetation | Tree-covered areas with dense canopy. Deep red in CIR imagery due to high near-infrared reflectance. Uneven texture with large shadows. |
| 6 | Herbaceous/Low Vegetation | Short vegetation, including grasses and shrubs. Smoother texture than forest, lighter red in CIR imagery. Common in clearings and field edges. |
| 7 | Cultivated Crops | Actively farmed fields, often in distinct rows. Bright red or pink in CIR imagery at peak growth. Presence of rows or patterns are used to distinguish from herbaceous areas. |
| 8 | Unclassified | Shadows or occlusions obscuring surface features. Deep black regions near trees and buildings. |
@@
3.2 Self-supervised pre-training
The pre-training phase of the methodology leverages the large dataset (377,921) of unannotated samples such that a ResNet-101 encoder can be trained to extract relevant features from the imagery that are well-suited for downstream land cover classification tasks. We compared two prominent self-supervised learning approaches: BYOL and MoCoV2. BYOL uses a self-distillation approach to learn representations from images using two separate but interdependent networks: an online network containing the image encoder, a projection head, and a prediction head, and a target network that is structurally similar to the online network but lacks the prediction head, initialized using a copy of the online network’s parameters at the start of training [36]. Given an input image, two separate augmented views of the image are created using a set of transformations. As illustrated in Figure 5, each view is then passed through the online and target networks to produce feature vectors. The online network, which processes view , is then trained to predict the feature vector produced by the target network from view . However, the target network is not trained directly; instead, it is updated using an exponential moving average (EMA) of the online network weights, giving the online network a stable yet constantly updated target to learn from. Ultimately, the image encoder from the online network is retained after training to be transferred to downstream tasks.
@@ 3.2 Self-supervised pre-training
BYOL uses an asymmetrical set of augmentations: the base set of augmentations T consists of random crops (scale , ratio ), horizontal flips (50% probability), random color jitter (brightness = 0.4, contrast = 0.4, saturation = 0.2, hue = 0.2; 80% probability), random grayscale conversion (20% probability), and Gaussian blur (kernel size = 25, ). A separate set of augmentations is also used, containing mostly the same set of augmentations as ) but with an additional random solarization procedure (threshold=0.5, 20% probability) and reducing the probability of Gaussian blur being applied to 10% from 100%. This asymmetry is enforced during training: given an image, each network will receive a view with augmentations drawn from either or , but both networks will never receive a view augmented using the same set of augmentations (e.g., both the online and target networks can not receive two randomly augmented images with augmentations in ).
Unlike BYOL, MoCoV2 is a contrastive learning paradigm that trains a network to maximize the cosine similarity between positive pairs (i.e., pairs of embeddings produced by a network given two random augmentations of the same image) while simultaneously minimizing the cosine similarity between negative pairs (i.e., pairs of embeddings derived from augmentations of different images) [43, 20]. MoCoV2 is closely related to SimCLR, a straightforward contrastive pretext task where the set of positive and negative pairs used to calculate loss is derived from a batch of images passed through an encoder network. While straightforward, contrastive learning works best with large numbers of negative samples, and thus SimCLR (and similar contrastive pre-training techniques) requires large in-memory batches for best performance [18, 15]. MoCoV2 addresses this by using a queue of negative pairs for the contrastive loss: embeddings are placed in the queue on the fly during training and serve as negative samples until they are eventually removed from the queue. This enables practical contrastive learning with large amounts of negative pairs, as the queue only needs to hold embeddings, not full images. To ensure the embeddings in the queue remain consistent despite the ever-evolving nature of the encoder network, a momentum (key) encoder network is used to construct the embeddings that will be enqueued. Like the teacher network in BYOL, this momentum encoder network is simply a slow moving average of the primary (query) encoder that is being trained. Unlike BYOL, these networks are symmetrical: each consists of a convolutional backbone (in our case ResNet-101) and a projection head. Like BYOL, however, the query and key encoders receive different augmentations of the same image. The loss is calculated according to the InfoNCE loss shown in Equation 3, where is the embedding produced by the query encoder, is the embedding produced by the key encoder (the positive pair), is the queue of negative samples, and is the temperature parameter (set to 0.2 per the original MoCoV2 implementation).
@@
The InfoNCE loss is backpropagated over the query encoder’s parameters, while the key encoder is updated according to Equation 2, where , corresponds to the key encoder, and corresponds to the query encoder. This momentum term is fixed throughout training to match the original MoCoV2 implementation. is then immediately placed in the queue of negative samples; if the queue size exceeds 65,536, the oldest embeddings in the queue are dequeued and discarded. After training is finished, the backbone of the query encoder is transferred to the downstream task, while the projection head and key encoder are discarded. MoCoV2 draws its set of augmentations from SimCLR, with random crops (scale , ratio ), random horizontal flips (50% probability), random color jitter (brightness = 0.4, contrast = 0.4, saturation = 0.4, hue = 0.1; 80% probability), random grayscale conversion (20% probability), and random Gaussian blur (kernel size = 25, ; 20% probability) being used to generate augmented views of images. We applied both BYOL and MoCoV2 pre-training to a ResNet-101 encoder network using the dataset of 377,921 unannotated samples from the state of Mississippi. We tested two configurations of each pre-training process: one starting with ImageNet pre-trained weights at the beginning of the pre-training process and one starting with randomly initialized weights. The pre-training process was run for 300 epochs, with the learning rate reduced from to 0 over the course of training using a cosine annealing schedule, with a warmup period of 10 epochs to linearly increase the learning rate from 0 to . The AdamW optimizer was used with a batch size of 4096: to achieve this large batch size while working within the memory constraints of a single Nvidia A100 GPU with 80 GB of VRAM, we employed gradient accumulation using a microbatch size of 256. After each epoch, the model was evaluated on the pre-training validation dataset using the pretext task’s loss function (as during training) to monitor for overfitting.
@@
To evaluate the quality of the representations learned during pre-training, we used a linear probing approach to determine how well the deep features extracted by the pre-trained networks could be used to classify land cover classes using a simple linear classifier. To accomplish this, the feature maps produced by the ResNet-101 encoder were upsampled to the original input resolution using bilinear interpolation, and a linear classifier was trained on top of these feature maps using the annotated training dataset from Section 3.1. The linear classifier consisted of a single convolutional layer with a softmax activation function to produce class probabilities for each pixel. K-fold cross-validation was employed to minimize the influence of variation in the training dataset while evaluating the performance of this method across different training dataset sizes ; the splitting strategy for each is shown in Figure 6.
@@
A warmup period of 10 epochs was used to linearly increase the learning rate from to at epoch 10, after which the learning rate was decayed by a factor of 0.1 when the validation loss failed to decrease for 10 epochs. After 50 epochs of no improvement in the validation loss, training was halted and the model with the lowest validation loss was selected as the final model for that run. Automatic mixed precision (AMP) training was used to speed up the training stage and reduce memory usage. Due to the limited size of the training datasets, we employed a strong set of image augmentations on-the-fly to ensure that the models did not overfit to the training data, including random horizontal and vertical flips (50% probability each), random 90∘ rotations (0∘, 90∘, 180∘, 270∘), color jitter (brightness = 0.4, contrast = 0.4, saturation = 0.2, hue = 0.2; 80% probability), random grayscale conversion (20% probability), and Gaussian blur (kernel size = 25, ). Since each cross-validation run used its validation split to tune the optimizer and trigger early stopping, we did not use validation-split metrics for final performance reporting. Rather, the performance of the linear probing models was evaluated using the point-based assessment dataset (25,000 samples), with the results being compared between the two MoCoV2 and two BYOL pre-training configurations (i.e., different weight initialization strategies) as well as a baseline model that used the ResNet-101 encoder with ImageNet pre-trained weights without any further pre-training. Considering the natural imbalance of the land cover classes, we chose the macro F1 score as the primary evaluation metric, as shown in Equation 5, where is the number of classes, is the number of true positives for class , is the number of false positives for class , and is the number of false negatives for class .
@@
Once the ResNet-101 encoders were pre-trained, they were used as a backbone for a semantic segmentation architecture that was fine-tuned using the annotated training dataset from Section 3.1. We evaluated 6 different semantic segmentation architectures to determine which framework was best suited for the land cover classification task at hand. We excluded any architectures that rely on transformer-based encoders (e.g., SegFormer [140], SETR [152]) or specialized convolutional backbones (e.g., HRNet [129], BiSeNetV2 [147]) to maintain a controlled comparison between architectures that all use the same ResNet-101 encoder network. Table 2 shows the semantic segmentation architectures that were evaluated in this study, and a brief description of each architecture is provided below.
| Model | Number of Parameters | Reference |
| FCN | 42,528,856 | Long et al. [76] |
| U-Net | 92,302,024 | Ronneberger et al. [103] |
| Attention U-Net | 83,610,272 | Oktay et al. [96] |
| DeepLabV3+ | 58,199,128 | Chen et al. [17] |
| UPerNet | 73,251,400 | Xiao et al. [139] |
| PAN | 140,688,970 | Li et al. [64] |
@@
U-Net is a popular semantic segmentation architecture that was among the early adopters of the encoder-decoder framework for semantic segmentation tasks. Under such a framework, the encoder network extracts hierarchical features from the input image, while the decoder network learns a mapping from these low-dimensional features to an output segmentation map that matches the original input resolution [103]. Notably, U-Net’s decoder network is approximately symmetric to the encoder network, meaning that the feature maps are upsampled in stages that mirror the downsampling stages of the encoder. As such, skip connections can be used to pass feature maps from the encoder to the decoder at corresponding stages: the encoder’s feature maps are concatenated with the decoder’s feature maps at each stage to provide the decoder with relevant, resolution-specific information that otherwise would be lost during the downsampling process. U-Net’s simple yet powerful architecture has made it a popular choice for a wide variety of remote sensing applications [102, 146, 65, 108] Our U-Net implementation differed from the original U-Net architecture in two key ways: first, we opted to use simple bilinear interpolation for upsampling in the decoder rather than transposed convolutions to reduce the presence of checkerboard artifacts in the output segmentation maps [95]; second, as we used a ResNet-101 encoder that downsamples the input image by a factor of 4 in the first stage (as opposed to the factor of 2 downsampling in the original U-Net), we modified the decoder to account for this difference by adding an additional upsampling stage to ensure that the output segmentation map matches the input resolution.
@@
DeepLabV3+ is an evolution of the DeepLab family of semantic segmentation architectures that incorporates atrous (dilated) convolutions and spatial pyramid pooling to capture multi-scale context in the feature maps [17]. At the core of DeepLabV3+ is an Atrous Spatial Pyramid Pooling (ASPP) module that aggregates multi-scale context from the final feature maps produced by the encoder. This ASPP module uses parallel atrous convolutions to encode feature maps at multiple scales as opposed to using pooling operations: the use of atrous (i.e., dilated) convolutions allows the network to capture features at multiple scales in a learned manner. A comparatively simple decoder network is utilized that upsamples the output of the ASPP module using bilinear interpolation, concatenates it with low-level feature maps from the encoder, and upsamples the result to produce the final segmentation map. Like U-Net, DeepLabV3+ has seen wide adoption for remote sensing applications [30, 72, 61]. DeepLabV3+ was originally designed to use a modified ResNet-101 encoder, and as such minimal effort was necessary to replace the original encoder with our pre-trained ResNet-101 encoder.